If a coin is tossed 5 times what is the probability it will land on heads?
There is a 50% chance that it will land on heads each toss. You need to clarify the question: do you mean what is the probability that it will land on heads at least once, exactly once, all five times?
What is the probability of a flipped coin landing on its side?
The probability of a coin landing head-side down is 0.5
The probability of landing head-side up is 0.5
Did you mean to ask the probability of it landing on its edge ?
How do I Find the mean of negative numbers?
You find the mean of negative numbers the same way you find the mean of any numbers. You add them up and then you divide by the number of numbers.
An example, using -3, -4, -6, and -9. The sum is -22. N is 4. The mean is -5.5.
The probability of flipping a coin twice and getting two heads is?
The probability of this is 50%. 2/4
True.
The probability of rolling a number greater than 5 on one roll of a die?
Only 6 is greater than 5 on a standard cubic die. The chance of rolling a 6 is one in six.
Two dice are thrown what are the odds in favor of rolling a score of six?
The first die can come up in any one of 6 ways.
The second die can come up in any one of 6 ways.
There are (6 x 6) = 36 different ways that two dice can fall.
There are 5 ways that they can come up as a six:
1 + 5
5 + 1
2 + 4
4 + 2
3 + 3
The probability of rolling a 6 is (5/36) = 13.89 %
The odds in favor are 5 to 31.
1.96
Tackling multiple choice questions relies primarily on the memory process called?
Tackling multiple choice questions relies primarily on the memory process called recognizing.
George Wickham did not attend the ball at Netherfield, because he did not want to meet Mr. Darcy
From a deck of 52 cards one card is drawn what is the probability that the card will be a red ten?
Since there are 2 red tens in a 52 card deck, probability of drawing one of the 2 red tens is 2/52 or 1/26 or 0.0385.
What are chances of getting heads 24 times in a row?
The chances of getting heads 24 times in a row is 1 in 224, or 1 in 16,777,216.
There are two possible sums that add up to a number evenly divisible by 2 and 3, namely 6 and 12.
There are 5 different ways to roll a sum of 6, so the probability of that happening is 5/36.
There is 1 way to roll a sum of 12, so the probability of that happening is 1/36.
The probability of either happening is the sum of the two separate probabilities.
p=5/36+1/36=6/36=1/6 ~= .1667 = 16.67%
What two unprecedented events occurred in connection with the ERA voting?
The first event was that when the ERA did not get ratified in its seven year limit, Congress extended ratification time by three extra years. The second event was that even though the deadline was extended the bill did not get ratified, and five states even tried to take back their votes.
Systematic sampling
What does simple random sample mean in statistics?
It means that every member of the population has the same probability of being included in the sample.
What is the probability that the sum of the faces is 7?
The probability that the sum of two dice is 7 is 6 in 36, or 1 in 6.
Of all the combinations, this is the one with the highest probability.
It is 6.1, approx.
A dice is thrown 3 times then sample space?
(1,1,1) (1,1,2) (1,1,3) (1,1,4) (1,1,5) (1,1,6) (1,2,1) (1,2,2) (1,2,3) (1,2,4) (1,2,5) (1,2,6) (1,3,1) (1,3,2) (1,3,3) (1,3,4) (1,3,5) (1,3,6) (1,4,1) (1,4,2) (1,4,3) (1,4,4) (1,4,5) (1,4,6) (1,5,1) (1,5,2) (1,5,3) (1,5,4) (1,5,5) (1,5,6) (1,6,1) (1,6,2) (1,6,3) (1,6,4) (1,6,5) (1,6,6) (2,1,1) (2,1,2) (2,1,3) (2,1,4) (2,1,5) (2,1,6) (2,2,1) (2,2,2) (2,2,3) (2,2,4) (2,2,5) (2,2,6) (2,3,1) (2,3,2) (2,3,3) (2,3,4) (2,3,5) (2,3,6) (2,4,1) (2,4,2) (2,4,3) (2,4,4) (2,4,5) (2,4,6) (2,5,1) (2,5,2) (2,5,3) (2,5,4) (2,5,5) (2,5,6) (2,6,1) (2,6,2) (2,6,3) (2,6,4) (2,6,5) (2,6,6) (3,1,1) (3,1,2) (3,1,3) (3,1,4) (3,1,5) (3,1,6) (3,2,1) (3,2,2) (3,2,3) (3,2,4) (3,2,5) (3,2,6) (3,3,1) (3,3,2) (3,3,3) (3,3,4) (3,3,5) (3,3,6) (3,4,1) (3,4,2) (3,4,3) (3,4,4) (3,4,5) (3,4,6) (3,5,1) (3,5,2) (3,5,3) (3,5,4) (3,5,5) (3,5,6) (3,6,1) (3,6,2) (3,6,3) (3,6,4) (3,6,5) (3,6,6) (4,1,1) (4,1,2) (4,1,3) (4,1,4) (4,1,5) (4,1,6) (4,2,1) (4,2,2) (4,2,3) (4,2,4) (4,2,5) (4,2,6) (4,3,1) (4,3,2) (4,3,3) (4,3,4) (4,3,5) (4,3,6) (4,4,1) (4,4,2) (4,4,3) (4,4,4) (4,4,5) (4,4,6) (4,5,1) (4,5,2) (4,5,3) (4,5,4) (4,5,5) (4,5,6) (4,6,1) (4,6,2) (4,6,3) (4,6,4) (4,6,5) (4,6,6) (5,1,1) (5,1,2) (5,1,3) (5,1,4) (5,1,5) (5,1,6) (5,2,1) (5,2,2) (5,2,3) (5,2,4) (5,2,5) (5,2,6) (5,3,1) (5,3,2) (5,3,3) (5,3,4) (5,3,5) (5,3,6) (5,4,1) (5,4,2) (5,4,3) (5,4,4) (5,4,5) (5,4,6) (5,5,1) (5,5,2) (5,5,3) (5,5,4) (5,5,5) (5,5,6) (5,6,1) (5,6,2) (5,6,3) (5,6,4) (5,6,5) (5,6,6) (6,1,1) (6,1,2) (6,1,3) (6,1,4) (6,1,5) (6,1,6) (6,2,1) (6,2,2) (6,2,3) (6,2,4) (6,2,5) (6,2,6) (6,3,1) (6,3,2) (6,3,3) (6,3,4) (6,3,5) (6,3,6) (6,4,1) (6,4,2) (6,4,3) (6,4,4) (6,4,5) (6,4,6) (6,5,1) (6,5,2) (6,5,3) (6,5,4) (6,5,5) (6,5,6) (6,6,1) (6,6,2) (6,6,3) (6,6,4) (6,6,5) (6,6,6)
What are the odds of getting 6 tails on 6 coin tosses?
If the coin is fair, then
The chance of tails on the 1st toss is 50% (1/2).
The chance of tails on the 2nd toss is 50% (1/2).
The chance of tails on the 3rd toss is 50% (1/2).
The chance of tails on the 4th toss is 50% (1/2).
The chance of tails on the 5th toss is 50% (1/2).
The chance of tails on the 6th toss is 50% (1/2).
The probability of getting all of them is (1/2) x (1/2) x (1/2) x (1/2) x (1/2) x (1/2) = 1/64 = 1.5625 % .
The odds are 63 to 1 against.
7/(4+7+5) = 7/16 = 43.75%
The mean is 12 and each observation is 8 units away from 12.
1 out of 15
Probab. = Prob. of red x Prob. of blue
Probab. = (3/10)x(2/9) = 5/90 = 1/15